### v1

@@ VI-A Experimental Setup
Fig. 10 shows the crack filling robot prototype, and Fig. 10 shows the indoor experimental setup. Instead of building actual cracks on the floor surface, we drew crack maps on drop cloths with blue paint to mainly test and validate the motion planner and robot control design. The robot delivered red paint to cover the drawn cracks to emulate the crack-filling action. The paint we selected for the experiments was dense, so it will not be dispensed or enlarged much after dropping on the cloth. By using such an arrangement, we can test the proposed motion planners and compare their performance with other benchmark algorithms under various crack characteristics (e.g., distribution and density) at a relatively low cost.
@@
Using optical markers placed on the top surface of the robot and the motion capture systems (8 Vantage cameras, Vicon Ltd.), we obtained the position and orientation of the robot at an updating frequency of $100$ Hz. A hydraulic pump and a solenoid valve were used for fluid paint delivery through the nozzle. With the known robot location, the local crack images within the region centered around the robot with a range of $S$ were fed to the planner to emulate the onboard crack detection sensor. The local position of the nozzle in the robot frame was obtained from a stereo camera mounted at the center of the robot pointing downward; see Fig. 10. In addition to the nozzle position, the stereo camera also provided the real-time location of the unfilled and filled cracks within $\mathcal{F}$.
@@
The onboard control implementation consisted of two platforms: the low-level controller was implemented on a real-time embedded system (Compact RIO NI-cRIO-9074, National Instruments Inc.), and the upper-level controller was a portable high-performance micro-processor (Intel NUC7i7DNK, Intel Corp.). The low-level controller mainly dealt with robot motion control and crack detection and planning. The imaging processing to identify the crack positions and the stepper motor control for nozzle motion were performed on the upper-level controller. The imaging data collection and motion control were implemented at $10$ Hz. Synchronization between the motion capture system and the onboard computers was implemented through a WiFi wireless connection.
@@
As shown in Fig. 10, the size of the indoor testing site is $l=5.79$ m and $w=6.10$ m. The physical and model parameters for the robot are: $R_{w}=7.6$ cm, $R_{g}=1.3$ cm, $R_{d}=49$ cm, $S=69$ cm, and $a=8.9$ cm. For the crack-filling planning and MPC design, we have the following parameters: $L=3/5a$, $\Delta=0.1$ s, $H=10$, $\omega_{r}^{\max}=1.31$ rad/s, $\omega_{n}^{\max}=7.87$ rad/s, and $v_{m}=0.1$ m/s. To create the crack maps with different densities and distributions, we used a crack image database in [44]. The density of a crack represents the Minkowski sum area of each topology (calculated by $\mathcal{T}\oplus S$) over the total workspace area. Random coordinate points and angles were generated by uniform and Gaussian distributions for the location and orientation of each topology. Four sets of crack maps were selected with different crack distributions and densities, that is, uniformly distributed cracks with a 100% density (denoted as U100) and an 80% density (U80), and Gaussian distribution cracks with a 100% density (G100) and a 20% density (G20).
@@
Figure 11: Experimental results for the robot and nozzle planning and control to fill a crack map that was generated by 80% density with uniform distribution (U80). (a) Trajectories of the robot and the nozzle. The robot’s center path is represented by the dashed black curve, and the nozzle path is marked by the solid orange line. The thick orange lines indicate the painted areas. The arrows dictate the robot’s traveling direction. The orange marks “ $\blacktriangle$” and “ $\bullet$” indicate the robot’s starting and ending locations. (b) A zoomed-in image of the robot and nozzle path for the portion marked in the blue rectangular box in (a). The thick light-blue and orange lines indicate the cracks and painted areas, respectively. The thick and thin red line segments indicate the active (i.e., delivering paint) and inactive filling (i.e., no paint delivery) actions, respectively. The red marks “ $\blacktriangle$” and “ $\bullet$” indicate the nozzle’s starting and ending locations for the zoomed-in path, respectively. Similarly, the blue marks “ $\blacksquare$” and “ $\bullet$” indicate the robot’s starting and ending locations for the zoomed-in path, respectively. (c) A plot of the nozzle trajectory with respect to the robot frame for the highlighted trajectory (from points $N_{0}$ to $N_{7}$) in (b). The red dotted circular is region $\mathcal{F}$. The thick and thin line segments indicate the active and inactive filling actions. The blue mark “ $\blacktriangle$” is the starting location of the nozzle, while red the “ $\bullet$” indicates the nozzle’s ending location.
@@ VI-B Evaluation Metrics
For comparison purposes, we also implemented two heuristic benchmark coverage planning algorithms. The first algorithm, denoted as ${\tt ZigZag}$, was to generate zigzag paths using the robot footprint coverage [26]. The ${\tt ZigZag}$ method generated zigzag motion with a slice width of $2a$ (i.e., diameter of $\mathcal{F}$) to solve the complete footprint coverage problem. The robot followed the zigzag path and only filled the cracks under its footprint area along the path. The ${\tt ZigZag}$ algorithm presented exhaustive coverage using the robot footprint, but the coverage contained large overlap. A greedy algorithm, denoted as ${\tt Greedy}$, was used to generate zigzag waypoints that cover the free space using the onboard detection sensor. If any targets are detected, the robot follows and covers them by footprint in the explored slices, then returns to the next unexplored waypoint to continue scanning for the targets. Algorithm 6 illustrates the implementation of the ${\tt Greedy}$ planner.
@@
Output : Path $\mathcal{P}_{R}$
1 $\mathcal{P}_{R}\leftarrow{\tt complete\_coverage}(\mathcal{W},S)$ for each unexplored waypoint of $\mathcal{P}_{R}$ $\rightarrow{P_{\text{n}}}(i)$ do
      2 ${\tt follow\_path}(P_{\text{n}}(i))$, $\mathcal{T}\leftarrow{\tt get\_topology}(\mathcal{I})$ if crack is found then follow $\mathcal{T}$
@@
We use several metrics to evaluate the motion planning and crack-filling performance: (1) filling time: the total time for the nozzle to deliver paint for all cracks; (2) robot traveling time: the total time that the robot takes to scan and fill the cracks; (3) robot path length and nozzle path length: the total arc lengths of the robot and the nozzle traveling trajectories, respectively; (4) sensor coverage: the percentage of the total sensor-covered areas over the entire work space. Sensor coverage values can be greater than 100% and it quantifies the overlapped coverage ratio; and (5) filling accuracy: the percentage fraction of the total length of cracks with filling error that is greater than a threshold. As shown in Fig. 10, we calculate the crack filling error as the difference between the extracted center lines of the cracks (blue line) and the filling paint (red line). The threshold value is taken as the variation of the width of the painted marks along the crack. In this study, we calculated and used $5$ mm as the threshold value.
@@ VII Results
VII-A Experimental Results
We first present the experimental results under the ${\tt oSCC}$ planner and nozzle motion control. Fig. 11 shows the robot and nozzle motion trajectory to fill a crack map that was generated by an 80% density with a uniform distribution (i.e., U80). Under the ${\tt oSCC}$ planner, the robot started from the left-upper corner and covered the entire workspace (as the black dash lines). The nozzle trajectory was shown as the orange lines to fill all the cracks. Fig. 11 further shows the robot and nozzle motion trajectories when first encountering multiple cracks within $\mathcal{F}$, that is, the zoom-in view of the blue boxed area shown in Fig. 11. The cracks and filled paint were marked by the light black and orange areas, respectively. At position $N_{0}$, the nozzle was at the center of the robot and traveled to position $N_{1}$ where filling of the crack commenced. The nozzle completed the segment between points $N_{1}$ and $N_{2}$. At $N_{2}$, the nozzle motion planner worked to switch among two cracks: it filled the segment between $N_{2}$ and $N_{3}$ before switching to complete the segment between $N_{2}$ and $N_{4}$. Filling stopped at $N_{4}$ and the nozzle then moved to position $N_{5}$, where the segment connecting to $N_{6}$ was filled. At $N_{6}$ the nozzle completed the filling and, through the motion of the robot, moved to point $N_{7}$. Other portions of the cracks in Fig. 11 that were not covered by nozzle motion from $N_{0}$ to $N_{7}$ were filled by the robot’s return trajectory (as shown in Fig. 11).
Fig. 11 plots the nozzle motion trajectory viewed in robot body frame $\mathcal{B}$ from $N_{0}$ to $N_{7}$. It is clear that most of the trajectories are within $\mathcal{F}$ (marked as the red dotted circle). Note, we consider the inner circle of the rectangle robot as the footprint $\mathcal{F}$ in the planners. The nozzle used in the experiment has the capacity to extend beyond $\mathcal{F}$ to fill the cracks. Fig. 12 shows the robot and nozzle velocity magnitude profiles for the experiment in Fig. 11. The plots include the robot’s traveling velocity magnitude (black line), the nozzle relative velocity magnitude profiles during filling cracks (thick orange dots), and no filling (orange dash lines). These results confirm the velocity constraint in (V-Cd) that the nozzle traveling velocity during filling actuation was much larger than the robot travel velocity, while the robot moved faster when the nozzle was not in filling action. These results clearly demonstrate how the nozzle planner works together with the robot motion.
We next present the comparison results among the various coverage planning algorithms. The experiments were conducted on four crack patterns: U100, U80, G100, and G20, with four planners: ${\tt oSCC}$, ${\tt SCC}$, ${\tt Greedy}$, and ${\tt ZigZag}$. Fig. 13 shows the experimental comparison of actual crack filling outcomes. Readers can further refer to the companion video clip for experimental comparisons under the four algorithms. By experimental outcomes, all of the planners completely covered the free space and filled the cracks. The planned trajectories under ${\tt oSCC}$ and ${\tt SCC}$ were different due to the online feature of the former algorithm. The trajectories under ${\tt Greedy}$ and ${\tt ZigZag}$ shared a similar zigzag scanning pattern, but the latter generated a much denser pattern because of the smaller size of the footprint $\mathcal{F}$ than the sensor coverage range (i.e., $a<S$).
@@ VII-A Experimental Results
Figure 14: Computational time comparisons of the ${\tt oSCC}$, ${\tt SCC}$, ${\tt GCC}$, ${\tt Greedy}$, and ${\tt ZigZag}$ planning algorithms for uniformly distributed cracks and Gaussian-based distributed cracks with different densities in the free space. (a) Total computation time to generate paths for uniformly distributed cracks. (b) Total computation time to generate paths for cracks with a Gaussian distribution. (c) Average computation time per iteration for cracks with uniform distribution. (d) Average computation time per iteration for Gaussian-based distributed cracks. The lines represent the mean values, while the shaded areas represent one standard deviation.
@@
We further conducted image processing to compute the evaluation metrics and compare the results under these planners. Table I lists the performance comparison. The ${\tt GCC}$ algorithm generates the optimal robot path for the footprint coverage problem. When the sensor range exceeds half the length of the rectangular workspace, the sensor can detect all cracks within the free space in a single scan, effectively reducing the SIFC problem to a footprint coverage problem. Therefore, the path lengths obtained from the ${\tt GCC}$ planner represent the lower bounds on the path length. For comparison purposes, the ${\tt GCC}$ planner was also implemented and included in the table as the benchmark. In Table I, the numbers in parentheses show the simulation results under the corresponding planner. We included these simulation results to validate the computational approach. We will present additional computational results later in this section. From the evaluation metrics in Table I, we observe several facts. First, the ${\tt oSCC}$ and ${\tt SCC}$ outperformed the ${\tt Greedy}$ and ${\tt ZigZag}$ planners in terms of filling time, robot traveling time, robot path length, nozzle path length, and sensor coverage. The filling accuracy values under all planners are similar and close to 100%. Between ${\tt oSCC}$ and ${\tt SCC}$, the performance is similar in terms of all evaluation metrics, and both are near-optimal compared with the ${\tt GCC}$ optimal planner. The sensor coverage values under all planners except ${\tt GCC}$ are more than 100%, which indicates the overlapped coverage under most planners. These comparison results confirmed that the efficiency and effectiveness of the ${\tt oSCC}$ planner achieved complete and near-optimal sensor and footprint coverage of the free space.
@@
To further evaluate the algorithms, we ran simulations to obtain the statistical performance of the motion planners with different crack densities and distributions. The dimensions of the workspace and all the parameters in the simulation setup are the same as the experimental ones. By randomly selecting topology and setting centroids and the orientation of the crack points and branch angles, five sets of uniformly distributed crack maps were generated. Each set consisted of eight maps with crack densities ranging from $35\%$ to $100\%$, that is, a total of $5\times 8=40$ uniformly distributed crack maps. Similarly, six crack maps per destiny were created by generating the coordinate points with a Gaussian distribution. The center of the map was the mean, and the area in percentage was the standard deviation. To achieve a different spread of the cracks, the standard deviations were set to $\frac{1}{12}$, $\frac{1}{6}$, and $\frac{1}{3}$ of the total workspace areas. This generated a total of $6\times 3\times 8=144$ Gaussian-distributed crack maps. Therefore, a total of 184 crack maps were generated and used for testing.
@@
We simulated and computed the performance of the ${\tt oSCC}$, ${\tt SCC}$, ${\tt GCC}$, ${\tt Greedy}$, and ${\tt ZigZag}$ planners with respect to different crack distributions (uniform and Gaussian) and densities. Figs. 14 and 14 show the total computation time comparison for the uniform and Gaussian distributions, respectively. The ${\tt oSCC}$ planner scanned and computed paths on an iterative basis, while others computed paths over the total workspace. In every iteration of the ${\tt oSCC}$, the robot scanned the area, extracted the crack graph, and computed the shortest path. The computation time of the ${\tt oSCC}$ depended on the crack density in the workspace. From the figures, the total computation time of the ${\tt oSCC}$ planner increases linearly with crack density. The computation time difference between ${\tt oSCC}$ and ${\tt SCC}$ reflects the iterative nature of the ${\tt oSCC}$ algorithm. In terms of computation time per iteration, the ${\tt oSCC}$ matches the range of the ${\tt SCC}$ planner, as shown in Figs. 14 and 14. The ${\tt oSCC}$ has a much lower computation time per iteration than that of the ${\tt Greedy}$ planner and outperforms both the ${\tt SCC}$ and the ${\tt GCC}$ at high crack densities.
@@
We next compared the robot path lengths and the sensor coverage against crack density. Fig. 15 shows the comparison of the robot path length and the sensor coverage for uniformly and Gaussian distributed cracks with different densities. The path lengths under the ${\tt GCC}$ and ${\tt ZigZag}$ planners are the lower and upper bounds, respectively. The statistical comparison confirms that the proposed ${\tt oSCC}$ achieves similar performance as the ${\tt SCC}$ algorithm, and both outperform the ${\tt Greedy}$ and ${\tt ZigZag}$ benchmarks. For the sensor coverage values, the ${\tt oSCC}$, ${\tt SCC}$, ${\tt Greedy}$, and ${\tt ZigZag}$ planners achieve $100\%$ workspace coverage for all the maps. The values greater than 100% measure the overlapped sensor coverage against the entire free space. Sensor overlapping shares the same trend as the path length comparison for ${\tt oSCC}$, ${\tt SCC}$, ${\tt Greedy}$, and ${\tt ZigZag}$. The overlapping area and robot path length under the ${\tt Greedy}$ algorithm are always larger than those by the ${\tt oSCC}$. Compared to the ${\tt Greedy}$ algorithm, the ${\tt oSCC}$ planner reduces the sensor overlap by up to $62\%$ and shortens the robot’s path by up to $24\%$ for densely and scatteredly distributed cracks. For uniformly distributed cracks, the ${\tt oSCC}$ and ${\tt SCC}$ planners achieve similar path lengths and sensor range coverage. For Gaussian-based distributed cracks, the ${\tt oSCC}$ planner results in less sensor overlap and shorter path length than ${\tt SCC}$. Therefore, the ${\tt oSCC}$ outperformed the other planners to cover and scan the cracks under both distributions.
@@
Figure 15: Robot path length and sensor coverage comparison under the ${\tt oSCC}$, ${\tt SCC}$, ${\tt GCC}$, ${\tt Greedy}$, and ${\tt ZigZag}$ planners for (a) uniformly distributed cracks and (b) Gaussian distributed cracks with different densities. The left and right $y$-axis labels represent the robot path length and the sensor coverage, respectively. The orange and green dotted lines show the sensor coverage and path length for ${\tt GCC}$, respectively. All planners except ${\tt GCC}$ achieve 100% sensor coverage. The lines and shaded areas represent the mean values and one standard deviation, respectively.

### v2

@@ VI-A Experimental Setup
Fig. 10 depicts the crack filling robot prototype, while Fig. 10 showcases the indoor experimental setup. Instead of creating actual cracks on the floor surface, we simulated crack maps on drop cloths using blue paint. This approach allowed us to primarily test and validate the motion planner and robot control design. To emulate the crack-filling action, the robot dispensed red paint to cover the drawn cracks. The paint chosen for the experiments was dense, minimizing dispersion or enlargement after application on the cloth. This setup enables us to assess the proposed motion planners and compare their performance with other benchmark algorithms across various crack characteristics, such as distribution and density, at a relatively low cost.
@@
Using optical markers positioned on the top surface of the robot and the motion capture system (consisting of 8 Vantage cameras, Vicon Ltd.), we captured the robot’s position and orientation at a frequency of $100$ Hz. A hydraulic pump and a solenoid valve were used for fluid paint delivery through the nozzle. With the known robot location, the local crack images within the region centered around the robot with a range of $S$ were fed to the planner to emulate the onboard crack detection sensor. The local position of the nozzle in the robot frame was obtained from a stereo camera mounted at the center of the robot pointing downward; see Fig. 10. In addition to the nozzle position, the stereo camera provided the real-time location of both unfilled and filled cracks within $\mathcal{F}$.
@@
The onboard control implementation consisted of two platforms: the low-level controller was deployed on a real-time embedded system (Compact RIO NI-cRIO-9074, National Instruments Inc.), while the upper-level controller was a portable high-performance microprocessor (Intel NUC7i7DNK, Intel Corp.). The low-level controller primarily handled robot motion control and crack detection and planning. Imaging processing to identify crack positions and stepper motor control for nozzle motion were performed on the upper-level controller. Imaging data collection and motion control were executed at a rate of $10$ Hz. Synchronization between the motion capture system and the onboard computers was achieved through a WiFi wireless connection.
@@
As shown in Fig. 10, the dimensions of the indoor testing site are $l=5.79$ m and $w=6.10$ m. The physical and model parameters for the robot are as follows: $R_{w}=7.6$ cm, $R_{g}=1.3$ cm, $R_{d}=49$ cm, $S=69$ cm, and $a=8.9$ cm. For the crack-filling planning and MPC design, the parameters are set as follows: $L=3/5a$, $\Delta=0.1$ s, $H=10$, $\omega_{r}^{\max}=1.31$ rad/s, $\omega_{n}^{\max}=7.87$ rad/s, and $v_{m}=0.1$ m/s. To create crack maps with varying densities and distributions, we used a crack image database from [44]. The density of a crack represents the Minkowski sum area of each topology (calculated by $\mathcal{T}\oplus S$) over the total workspace area. Random coordinate points and angles were generated using uniform and Gaussian distributions for the location and orientation of each topology. Four sets of crack maps were selected with different crack distributions and densities, that is, uniformly distributed cracks with a 100% density (denoted as U100) and an 80% density (U80), and Gaussian distribution cracks with a 100% density (G100) and a 20% density (G20).
@@
Figure 11: Experimental results for the robot and nozzle planning and control to fill a crack map that was generated by 80% density with uniform distribution (U80). (a) Trajectories of the robot and the nozzle. The robot’s center path is represented by the dashed black curve, and the nozzle path is marked by the solid orange line. The thick orange lines indicate the painted areas. The arrows dictate the robot’s traveling direction. The orange marks “ $\blacktriangle$” and “ $\bullet$” indicate the robot’s starting and ending locations. (b) A zoomed-in image of the robot and nozzle path for the portion marked in the blue rectangular box in (a). The thick gray and orange regions represent the cracks and painted areas, respectively. The thick and thin red line segments indicate the active (i.e., delivering paint) and inactive filling (i.e., no paint delivery) actions, respectively. The red marks “ $\blacktriangle$” and “ $\bullet$” indicate the nozzle’s starting and ending locations for the zoomed-in path, respectively. Similarly, the blue marks “ $\blacksquare$” and “ $\bullet$” indicate the robot’s starting and ending locations for the zoomed-in path, respectively. (c) A plot of the nozzle trajectory with respect to the robot frame for the highlighted trajectory (from points $N_{0}$ to $N_{7}$) in (b). The red dotted circular is region $\mathcal{F}$. The thick and thin line segments indicate the active and inactive filling actions. The blue mark “ $\blacktriangle$” is the starting location of the nozzle, while red the “ $\bullet$” indicates the nozzle’s ending location.
@@ VI-B Evaluation Metrics
For comparison purposes, we also implemented two heuristic benchmark coverage planning algorithms. The first algorithm, ${\tt ZigZag}$, solves the complete footprint coverage problem [26] by generating zigzag paths with a slice width of $2a$ (i.e., the diameter of $\mathcal{F}$). The robot then follows this path, filling only the cracks within its footprint area along the trajectory. The ${\tt ZigZag}$ algorithm provides exhaustive coverage using the robot’s footprint, but the coverage contains large overlaps. A greedy algorithm, denoted as ${\tt Greedy}$, is used to generate zigzag waypoints that cover the free space using the onboard detection sensor. If any targets are detected, the robot follows and covers them by footprint in the explored slices, then returns to the next unexplored waypoint to continue scanning for the targets. Algorithm 6 illustrates the implementation of the ${\tt Greedy}$ planner.
@@
Output : $\mathcal{P}_{R}$
1 $\mathcal{P}_{R}\leftarrow{\tt complete\_coverage}(\mathcal{W},S)$
for each unexplored waypoint of $\mathcal{P}_{R}$ $\rightarrow{P_{\text{n}}}(i)$ do
      2 ${\tt follow\_path}(P_{\text{n}}(i))$, $\mathcal{T}\leftarrow{\tt get\_topology}(\mathcal{I})$
       if crack is found then follow $\mathcal{T}$
@@ 2 ${\tt follow\_path}(P_{\text{n}}(i))$, $\mathcal{T}\leftarrow{\tt get\_topology}(\mathcal{I})$
We used various metrics to assess the performance of motion planning and crack filling: (1) filling time: the total time for the nozzle to deliver paint for all cracks; (2) robot traveling time: the overall time taken by the robot to scan and fill the cracks; (3) robot path length and nozzle path length: the combined arc lengths of the trajectories traveled by the robot and the nozzle, respectively; (4) sensor coverage: the percentage of the total sensor-covered area over the entire workspace. Sensor coverage values can exceed 100%, indicating overlapping coverage ratios; and (5) filling accuracy: the percentage fraction of the total length of cracks with filling error that is greater than a threshold. As shown in Fig. 10, we calculated the crack filling error as the difference between the extracted center lines of the cracks (blue line) and the filling paint (red line). The threshold value was taken as the variation in the width of painted marks along the crack. In this study, we calculated and used $5$ mm as the threshold value.
@@ VII Results
VII-A Experimental Results
We first present the experimental results under the ${\tt oSCC}$ planner and nozzle motion control. Fig. 11 shows the robot and nozzle motion trajectory to fill a crack map that was generated by an 80% density with a uniform distribution (i.e., U80). Under the ${\tt oSCC}$ planner, the robot started from the upper-left corner and covered the entire workspace (as indicated by the black dashed lines). The nozzle trajectory, depicted by the orange lines, effectively filled all the cracks. Fig. 11 provides a close look at the robot and nozzle motion trajectories upon encountering multiple cracks within $\mathcal{F}$, offering a zoomed-in view of the area outlined in blue in Fig. 11. The cracks and filled paint are represented by the gray and orange areas, respectively. At position $N_{0}$, the nozzle was aligned with the robot’s center and proceeded towards position $N_{1}$ to commence crack filling. It then moved on to complete the segment between points $N_{1}$ and $N_{2}$. Upon reaching $N_{2}$, the nozzle motion planner facilitated a switch between two cracks: it filled the segment between $N_{2}$ and $N_{3}$ before transitioning to complete the segment between $N_{2}$ and $N_{4}$. Filling ceased at $N_{4}$, prompting the nozzle to proceed to position $N_{5}$, where it filled the segment connecting to $N_{6}$. At $N_{6}$, the nozzle finalized the filling process and, through the motion of the robot, moved to point $N_{7}$. Remaining portions of the cracks in Fig. 11, not covered by the nozzle’s motion from $N_{0}$ to $N_{7}$, were filled during the robot’s return trajectory (as illustrated in Fig.11).
Fig. 11 illustrates the nozzle motion trajectory viewed in the robot body frame $\mathcal{B}$ from $N_{0}$ to $N_{7}$. Most of the trajectories are contained within $\mathcal{F}$, marked as the red dotted circle. It’s worth noting that we consider the inner circle of the rectangular robot as the footprint $\mathcal{F}$ in the planners. The experimental nozzle has the capability to extend beyond $\mathcal{F}$ to fill the cracks. Fig. 12 illustrates the velocity magnitude profiles for both the robot and nozzle during the experiment shown in Fig. 11. The plots include the robot’s traveling velocity magnitude represented by the black line, the nozzle’s relative velocity magnitude profiles during crack filling (depicted by thick orange dots), and during non-filling periods (shown by orange dashed lines). These results validate the velocity constraint specified in (8d), which indicates that the nozzle’s traveling velocity during filling actions was significantly higher than the robot’s travel velocity. Conversely, the robot moved faster when the nozzle was not engaged in filling actions. These findings underscore the collaborative operation between the nozzle planner and the robot motion.
@@ VII-A Experimental Results
Figure 14: Computational time comparisons of the ${\tt oSCC}$, ${\tt SCC}$, ${\tt GCC}$, ${\tt Greedy}$, and ${\tt ZigZag}$ planning algorithms for uniformly distributed cracks and Gaussian-based distributed cracks with different densities in the free space. (a) Total computation time to generate paths for uniformly distributed cracks. (b) Total computation time to generate paths for cracks with Gaussian distributions. (c) Average computation time per iteration for cracks with uniform distribution. (d) Average computation time per iteration for Gaussian-based distributed cracks. The lines represent the mean values, while the shaded areas represent one standard deviation.
We next present the comparison results among the various coverage planning algorithms. The experiments were conducted on four crack patterns: U100, U80, G100, and G20, using four planners: ${\tt oSCC}$, ${\tt SCC}$, ${\tt Greedy}$, and ${\tt ZigZag}$. Fig. 13 shows the experimental comparison of actual crack filling outcomes. Readers can further refer to the companion video clip for experimental comparisons under the four algorithms. By experimental results, all of the planners completely covered the free space and filled the cracks. The planned trajectories under ${\tt oSCC}$ and ${\tt SCC}$ were different due to the online feature of the former algorithm. The trajectories under ${\tt Greedy}$ and ${\tt ZigZag}$ shared a similar zigzag scanning pattern, but the latter generated a much denser pattern because of the smaller size of the footprint $\mathcal{F}$ than the sensor coverage range (i.e., $a<S$).
@@
We further conducted image processing to compute the evaluation metrics and compare the results under these planners. Table I lists the performance comparison. The ${\tt GCC}$ algorithm generates the optimal robot path for the footprint coverage problem. When the sensor range exceeds half the length of the rectangular workspace, the sensor can detect all cracks within the free space in a single scan, effectively reducing the SIFC problem to a footprint coverage problem. Therefore, the path lengths obtained from the ${\tt GCC}$ planner represent the lower bounds on the path length. For comparison purposes, the ${\tt GCC}$ planner was also implemented and included in the table as the benchmark. In Table I, the bolded values indicate the best performance among all five planners, and the numbers in parentheses show the simulation results under the corresponding planner. We included these simulation results to validate the computational approach. We will present additional computational results later in this section. From the evaluation metrics in Table I, we observe several facts. First, the ${\tt oSCC}$ and ${\tt SCC}$ outperformed the ${\tt Greedy}$ and ${\tt ZigZag}$ planners in terms of filling time, robot traveling time, robot path length, nozzle path length, and sensor coverage. The filling accuracy values under all planners are similar and close to 100%. Between ${\tt oSCC}$ and ${\tt SCC}$, the performance is similar in terms of all evaluation metrics, and both are near-optimal compared with the ${\tt GCC}$ optimal planner. The sensor coverage values under all planners except ${\tt GCC}$ are more than 100%, which indicates the overlapped coverage under most planners. These comparison results confirm the efficiency and effectiveness of the ${\tt oSCC}$ planner in achieving complete and near-optimal sensor and footprint coverage of the free space.
@@
To further evaluate the algorithms, we conducted simulations to assess the statistical performance of the motion planners with different crack densities and distributions. The dimensions of the workspace and all parameters in the simulation setup were kept consistent with the experimental setup. By randomly selecting topologies and setting centroids and the orientation of the crack points and branch angles, five sets of uniformly distributed crack maps were generated. Each set consisted of eight maps with crack densities ranging from $35\%$ to $100\%$, resulting in a total of $5\times 8=40$ uniformly distributed crack maps. Similarly, six crack maps per density were created by generating coordinate points with a Gaussian distribution. The center of the map was represented by the mean of the distribution, and the area in percentage was the standard deviation. To achieve different spreads of the cracks, the standard deviations were set to $\frac{1}{12}$, $\frac{1}{6}$, and $\frac{1}{3}$ of the total workspace areas. This generated a total of $6\times 3\times 8=144$ Gaussian-distributed crack maps. Therefore, a total of 184 crack maps were generated and used for testing.
@@
We simulated and computed the performance of the ${\tt oSCC}$, ${\tt SCC}$, ${\tt GCC}$, ${\tt Greedy}$, and ${\tt ZigZag}$ planners with respect to different crack distributions (uniform and Gaussian) and densities. Figs. 14 and 14 show the total computation time comparison for the uniform and Gaussian distributions, respectively. The ${\tt oSCC}$ planner scanned and computed paths on an iterative basis, while others computed paths over the entire workspace. In each iteration of the ${\tt oSCC}$, the robot scanned the area, extracted the crack graph, and computed the shortest path. The computation time of the ${\tt oSCC}$ depended on the crack density in the workspace. From the figures, the total computation time of the ${\tt oSCC}$ planner increases linearly with crack density. The computation time difference between ${\tt oSCC}$ and ${\tt SCC}$ reflects the iterative nature of the ${\tt oSCC}$ algorithm. In terms of computation time per iteration, the ${\tt oSCC}$ matches the range of the ${\tt SCC}$ planner, as shown in Figs. 14 and 14. The ${\tt oSCC}$ has a much lower computation time per iteration than that of the ${\tt Greedy}$ planner and outperforms both the ${\tt SCC}$ and the ${\tt GCC}$ at high crack densities.
@@
We next compared the robot path lengths and sensor coverage against crack density. Fig. 15 illustrates the comparison of robot path length and sensor coverage for uniformly and Gaussian distributed cracks with varying densities. The path lengths under the ${\tt GCC}$ and ${\tt ZigZag}$ planners represent the lower and upper bounds, respectively. The statistical comparison confirms that the proposed ${\tt oSCC}$ achieves similar performance as the ${\tt SCC}$ algorithm, and both outperform the ${\tt Greedy}$ and ${\tt ZigZag}$ benchmarks. Regarding sensor coverage values, the ${\tt oSCC}$, ${\tt SCC}$, ${\tt Greedy}$, and ${\tt ZigZag}$ planners achieve $100\%$ workspace coverage for all the maps. The values greater than 100% measure the overlapped sensor coverage against the entire free space. Sensor overlapping shares the same trend as the path length comparison for ${\tt oSCC}$, ${\tt SCC}$, ${\tt Greedy}$, and ${\tt ZigZag}$. The overlapping area and robot path length under the ${\tt Greedy}$ algorithm are always larger than those by the ${\tt oSCC}$. Compared to the ${\tt Greedy}$ algorithm, the ${\tt oSCC}$ planner reduces the sensor overlap by up to $62\%$ and shortens the robot’s path by up to $24\%$ for densely and scatteredly distributed cracks. For uniformly distributed cracks, the ${\tt oSCC}$ and ${\tt SCC}$ planners achieve similar path lengths and sensor range coverage. In the case of Gaussian-based distributed cracks, the ${\tt oSCC}$ planner results in less sensor overlap and shorter path length than ${\tt SCC}$. Therefore, the ${\tt oSCC}$ outperforms the other planners in covering and scanning the cracks under both distributions.
@@
Figure 15: Robot path length and sensor coverage comparison under the ${\tt oSCC}$, ${\tt SCC}$, ${\tt GCC}$, ${\tt Greedy}$, and ${\tt ZigZag}$ planners for (a) uniformly distributed cracks and (b) Gaussian distributed cracks with different densities. The left and right $y$-axis labels represent the robot path length and sensor coverage, respectively. The orange and green dotted lines show the sensor coverage and path length for ${\tt GCC}$, respectively. All planners except ${\tt GCC}$ achieve 100% sensor coverage. The lines represent the mean values, and the shaded areas indicate one standard deviation.